# The quantum-refractive evolution of polarization states in pulsar emission

### v1

@@ 3 Conclusions and discussion
We have investigated the evolution of polarization states under the quantum refraction effects combined with the frequency dependence of pulsar emission. To this end, we have solved a system of evolution equations of the Stokes vector given by (1) (or by (23)-(25)) for three examples of RPPs at a fixed frequency for specific emissions (e.g., X-rays as in Sections 2.2.1 and 2.2.2). Our main results are presented in Figs. 3-6, from numerical solutions and in part from perturbative solutions. Also, we have approximated the birefringent vector by some models as in Fig. 7 to solve the evolution equations analytically, and obtained the results as presented in Fig. 8, from approximate analytical solutions. It is noteworthy that at a fixed frequency of emission the evolution of the Stokes vector largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, in which the magnetic field strength is a dominant factor: (i) fractionally oscillatory - adiabatic/monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviours. These features are shown by the numerical solutions in Figs. 3-6, and also confirmed by the approximate analytical solutions in Fig. 8.
@@
This study is centred on solving the evolution equations for polarization states (1), wherein the birefringent vector that contains all information about the quantum refraction effects, coupled to the frequency of pulsar emission, acts on the Stokes vector; the evolution results from the combination of the quantum refraction effects and the frequency dependence of the emission. This is a major difference from our previous work (Kim et al. (2024)), wherein the same effects are independent of the emission frequency; the work solely focuses on the quantum refraction effects on the propagation and polarization of a photon from pulsar emission, with no reference to other properties of the emission, such as the frequency. In this regard, it is worthwhile to draw comparison between the two quantities, the polarization vector and the Stokes vector, both of which are used to describe polarization states. The polarization vector is defined directly from the radiative electric field vector (i.e. the unit electric field vector), and it is parallel-transported along the the propagation vector; usually, we consider such two vectors orthogonal to each other and to the propagation vector to define an orthonormal basis consisting of the three vectors. In contrast, the Stokes vector is defined from Stokes parameters which are built out of the radiative electric field vector (Rybicki & Lightman (1991)). The representation of the Stokes vector is abstract in the sense that it is a vector defined on the Poincaré sphere. The Stokes vector is not parallel-transported along the the propagation vector, but can still be defined along the propagation vector as the two polarization vectors move along it; hence, it can be parametrised by $s$ to represent polarization states along the photon trajectory. However, the Stokes vector has a crucial advantage over the polarization vector in representing polarization states in some astrophysical studies like this: it can be directly estimated from polarimetric measurements and accommodate depolarization effects due to incomplete coherence and random processes during the photon propagation (Scully & Zubairy (1997)).
With regard to the adiabatic evolution condition as mentioned in Heyl & Shaviv (2000), Heyl et al. (2003), and Heyl & Caiazzo (2018), we have carefully examined our results presented in Figs. 3-5 to see what interpretations the condition leads to. Solving the condition $\left|k\mathbf{\hat{\Omega}}\left(\mathrm{d}\ln|k\mathbf{\hat{\Omega}}|/\mathrm{d}s\right)^{-1}\right|\gtrsim 0.05$ for $s$ (note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in the references above and that we have set the condition value to $0.05$ rather than $0.5$ as in the references) yields $s_{\mathrm{PL}1}\lesssim s\lesssim s_{\mathrm{PL}2}$, where $s_{\mathrm{PL}1[2]}$ refers to the lower [upper] bound for the ‘polarization limiting’ distance as measured from the emission point. Using this, we have checked out the following: (1) for $6.2\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.8\times 10^{6}\,\mathrm{cm}$ in Fig. 3, (2) for $3.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 6.2\times 10^{6}\,\mathrm{cm}$ in Fig. 4, (3) for $1.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.9\times 10^{7}\,\mathrm{cm}$ in Fig. 5, our Stokes vector evolves evidently; otherwise, it freezes.
@@
Acknowledgements
D.-H.K was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1I1A1A01054781). C.M.K. was supported by Ultrashort Quantum Beam Facility operation program (140011) through APRI, GIST and GIST Research Institute (GRI) grant funded by GIST. S.P.K. was also in part supported by National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2019R1I1A3A01063183).
Data Availability
The inclusion of a Data Availability Statement is a requirement for articles published in MNRAS. Data Availability Statements provide a standardised format for readers to understand the availability of data underlying the research results described in the article. The statement may refer to original data generated in the course of the study or to third-party data analysed in the article. The statement should describe and provide means of access, where possible, by linking to the data or providing the required accession numbers for the relevant databases or DOIs.
@@ Appendix A The classical Stokes vector
where $t_{\mathrm{ret}}\equiv t-r/c$ is the retarded time, $\boldsymbol{\xi}$ represents the particle’s trajectory, $\mathbf{n}$ is the propagation direction of the radiation, and an over-dot ${\dot{}}$ denotes differentiation with respect to $t$. In a suitably chosen Cartesian frame, by setting $\boldsymbol{\xi}\left(t_{\mathrm{ret}}\right)=\rho\left(\sin\left(\beta ct_{\mathrm{ret}}/\rho\right),0,\cos\left(\beta ct_{\mathrm{ret}}/\rho\right)\right)$, with $\rho$ being the radius of curvature of the particle’s trajectory, and $\mathbf{n}=\left(\cos\varphi,\sin\varphi,0\right)$, with $\varphi$ being the angle measured from the $x$-axis to the plane of the particle’s motion, we can construct a simple toy model for pulse profiles of pulsar curvature emission as described below (Kim & Trippe (2021)).
One can express Stokes parameters out of the radiation field (46), which describe polarization properties of the curvature radiation (Gil & Snakowski (1990)):
@@
where $\tilde{E}_{\parallel}$ and $\tilde{E}_{\perp}$ denote the components of the Fourier transform $\mathbf{\tilde{E}}\left(\omega\right)=\tilde{E}_{\parallel}\left(\omega\right)\mathbf{e}_{z}+\tilde{E}_{\perp}\left(\omega\right)\,\mathbf{e}_{y}=\int_{-\infty}^{\infty}\mathbf{E}\left(t\right)\exp\left(\mathrm{i}\omega t\right)\,\mathrm{d}t$, expressed in the Cartesian frame, and means the complex conjugate, and $\mathcal{E}_{\mathrm{o}}=q\beta/\left(2\sqrt{3}\pi^{2}r\rho\right)$, and $\delta\equiv\gamma^{-1}=\left(1-\beta^{2}\right)^{1/2}\ll 1$ is the half-angle of the beam emission, and $\mathrm{K}_{1/3}$ and $\mathrm{K}_{2/3}$ denote the modified Bessel functions of the second kind. With regard to the polarization state of the curvature radiation, $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization. These parameters can be plotted as functions of the phase angle $\varphi$, where $\varphi\leq\delta\ll 5^{\circ}$ usually, to simulate the pulse profiles of pulsar emission theoretically.
@@
In Fig. 9 is plotted the classical Stokes vector against the phase angle $\varphi$, where we have set, for example, $\delta\approx 10^{-8}$, $\rho\approx 10^{8}\,\mathrm{cm}$, $\beta\approx 1$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ to model pulse profiles of X-ray pulsar emission. Here the initial values for the Stokes vector, $\mathbf{S}\left(\varphi=0\right)=\left(1,0,0\right)$ and $\mathbf{S}\left(\varphi\approx-1.16\times 10^{-6}\,\mathrm{rad}\right)=\left(0.8,0,0.6\right)$, as in the examples given in Section 2.2.1, are marked by solid circles and solid boxes, respectively.
Figure 9: The classical Stokes vector $\mathbf{S}\left(\varphi\right)=\left(S_{1}\left(\varphi\right),S_{2}\left(\varphi\right),S_{3}\left(\varphi\right)\right)$ plotted against the phase angle $\varphi$. Its initial values $\mathbf{S}\left(\varphi=0\right)=\left(1,0,0\right)$ and $\mathbf{S}\left(\varphi\approx-1.16\times 10^{-6}\,\mathrm{rad}\right)=\left(0.8,0,0.6\right)$ are marked by solid circles and solid boxes, respectively.
@@ Appendix B Approximate analytical solutions to evolution equations
Substituting equation (30) into equations (23)-(25), the evolution equations can be reduced as follows:
@@
First, we solve equation (56) for $S_{3}\left(s\right)$, and then using this solution, obtain $S_{1}\left(s\right)$ and $S_{2}\left(s\right)$, by integrating equations (54) and (55), respectively:
@@
Then $S_{\mathrm{o}}$, $C$ and $\delta$ are determined by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with equations (57)-(59) evaluated at $s=0$.

### v3

@@ 4 Conclusions and discussion
We have investigated the evolution of polarization states of pulsar emission under the quantum refraction effects, combined with the dependence on the emission frequency, for both dipole and quadrudipole pulsar models. To this end, we have solved a system of evolution equations of the Stokes vector given by (1) (or by (23)-(25)) and (59) (or by (66)-(68)) in the dipole and quadrudipole cases, respectively, for three examples of RPPs at a fixed frequency for specific emissions (e.g., X-rays as in sections 2.2.1, 2.2.2 and 3.2.1). Our main results are presented by Figs. 3-5 and 10-12 in the dipole and quadrudipole cases, respectively, from numerical solutions and in part from perturbative solutions. Also, we have replaced the birefringent vector with some approximate models as in Fig. 7 to solve the evolution equations analytically in the dipole case, and obtained the results as presented by Fig. 8. It is noteworthy that at a fixed frequency of emission the evolution of the Stokes vector largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, in which the magnetic field strength is a dominant factor: (i) fractionally oscillatory - monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviors. These features are shown by the numerical solutions in Figs. 3-6, and also confirmed by the approximate analytical solutions in Fig. 8. In addition, we have examined how the aforementioned features change in the quadrudipole case.
@@
This study is centered on solving the evolution equations for polarization states (1), wherein the birefringent vector that contains all information about the quantum refraction effects, coupled to the frequency of pulsar emission, acts on the Stokes vector; the evolution results from the combination of the quantum refraction effects and the frequency dependence of the emission. This is a major difference from our previous work 10.1093/mnras/stae1304 , wherein the same effects have no connection to the emission frequency; the work solely focuses on the quantum refraction effects on the propagation and polarization vectors in pulsar emission, with no reference to other properties, such as the emission frequency. In this regard, it is worthwhile to draw comparison between the two quantities, the polarization vector and the Stokes vector, both of which are used to describe polarization states. The polarization vector is defined directly from the radiative electric field vector (i.e., the unit electric field vector), and it is parallel-transported along the the propagation vector; usually, we consider such two vectors orthogonal to each other and to the propagation vector to define an orthonormal basis consisting of the three vectors. In contrast, the Stokes vector is defined from Stokes parameters which are built out of the radiative electric field vector Rybicki:847173 . The representation of the Stokes vector is abstract in the sense that it is a vector defined on the Poincaré sphere. The Stokes vector is not parallel-transported along the the propagation vector, but can still be defined along the propagation vector as the two polarization vectors move along it; hence, it can be parameterized by $s$ to represent polarization states along the photon trajectory. However, the Stokes vector has a crucial advantage over the polarization vector in representing polarization states in some astrophysical studies like this: it can be directly estimated from polarimetric measurements and accommodate depolarization effects due to incomplete coherence and random processes during the photon propagation Scully1997 .
Our results in this study may be of some significance for the currently operating and planned X-ray space telescopes: Imaging X-ray Polarimetry Explorer (IXPE) IXPE , X-ray Polarimeter Satellite (XPoSat) XPoSat , the enhanced X-ray Timing and Polarimetry mission (eXTP) Santangelo_2019 and the Compton Telescope project Wadiasingh2019Magnetars . These telescopes measure the polarization of the X-rays from energetic compact objects such as magnetars and black holes to unveil their geometry and physical environment in detail. Several magnetars observed by IXPE have been estimated to have overcritical field strengths Taverna2022Polarized ; Turolla2024IXPE . Furthermore, a recent measurement of the X-ray polarization of the magnetar 4U 0142+61 has shown that the polarization degree and angle change as a function of X-ray energy, the interpretation of which has led to two competing scenarios about the X-ray emission of the magnetar Taverna2022Polarized ; Lai2023IXPE . In fact, a full analysis of the polarimetry data would require a physical model that comprehensively incorporates the properties regarding the polarization of surface emission, the photon propagation through magnetized plasmas, birefringence due to a magnetized quantum vacuum, and gravitational effects on photon propagation Taverna2015Polarization ; Caiazzo2022Probing .
In this study, we have focused on vacuum birefringence as it is one of the most significant phenomenological issues to be tested by the X-ray polarimetry in practice. When the magnetic field is sufficiently strong and slowly varying, the polarization states evolve due to vacuum birefringence; that is, the Stokes vector components change during the photon propagation within the so-called polarization-limiting radius Heyl2000 ; Heyl2003highenergy , which can be several to a couple dozen neutron-star radii, depending on the magnetic field strength at the surface of a neutron star and the emission frequency. However, beyond the polarization-limiting radius, the polarization states ‘freeze’, remaining the same until finally being observed through polarimetry Heyl2003highenergy ; Caiazzo2022Probing . The new features of the polarization evolution presented in our study, such as the three different oscillatory patterns of the Stokes vectors and the effects of the possible contributions from the multipolar components may all closely concern the observation through the X-ray polarimetry, and therefore should be taken into proper consideration for a more accurate model for pulsar emission.
@@
[…]
Acknowledgments D.-H.K was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1I1A1A01054781). C.M.K. was supported by Ultrashort Quantum Beam Facility operation program (140011) through APRI, GIST and GIST Research Institute (GRI) grant funded by GIST. S.P.K. was also in part supported by National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2019R1I1A3A01063183).
@@ Appendix A The classical Stokes vector
where $t_{\mathrm{ret}}\equiv t-r/c$ is the retarded time, $\boldsymbol{\xi}$ represents the particle’s trajectory, $\mathbf{n}$ is the propagation direction of the radiation, and an over-dot ${\dot{}}$ denotes differentiation with respect to $t$. In a suitably chosen Cartesian frame, by setting $\boldsymbol{\xi}\left(t_{\mathrm{ret}}\right)=\rho\left(\sin\left(\beta ct_{\mathrm{ret}}/\rho\right),0,\cos\left(\beta ct_{\mathrm{ret}}/\rho\right)\right)$, with $\rho$ being the radius of curvature of the particle’s trajectory, and $\mathbf{n}=\left(\cos\varphi,\sin\varphi,0\right)$, with $\varphi$ being the angle measured from the $x$-axis to the plane of the particle’s motion, we can construct a simple toy model for pulse profiles of pulsar curvature emission as described below Kim2021Generala .
One can express Stokes parameters out of the radiation field (69), which describe its polarization properties Gil1990Curvature :
@@
where $\tilde{E}_{\parallel}$ and $\tilde{E}_{\perp}$ denote the components of the Fourier transform $\mathbf{\tilde{E}}\left(\omega\right)=\tilde{E}_{\parallel}\left(\omega\right)\mathbf{e}_{z}+\tilde{E}_{\perp}\left(\omega\right)\,\mathbf{e}_{y}=\int_{-\infty}^{\infty}\mathbf{E}\left(t\right)\exp\left(\mathrm{i}\omega t\right)\,\mathrm{d}t$, expressed in the Cartesian frame, and means the complex conjugate, and $\mathcal{E}_{\mathrm{o}}=q\beta/\left(2\sqrt{3}\pi^{2}r\rho\right)$, and $\delta\equiv\gamma^{-1}=\left(1-\beta^{2}\right)^{1/2}\ll 1$ is the half-angle of the beam emission, and $\mathrm{K}_{1/3}$ and $\mathrm{K}_{2/3}$ denote the modified Bessel functions of the second kind. With regard to the polarization state of the radiation field, $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization. These parameters can be plotted as functions of the phase angle $\varphi$, where $\varphi\leq\delta\ll 5^{\circ}$ usually, to simulate the pulse profiles of pulsar emission theoretically.
@@
In Fig. 14 is plotted the classical Stokes vector against the phase angle $\varphi$, where we have set, for example, $\delta\approx 10^{-8}$, $\rho\approx 10^{8}\,\mathrm{cm}$, $\beta\approx 1$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ to model pulse profiles of X-ray pulsar emission. Here the initial values for the Stokes vector, $\mathbf{S}\left(\varphi=0\right)=\left(1,0,0\right)$ and $\mathbf{S}\left(\varphi\approx-1.16\times 10^{-6}\,\mathrm{rad}\right)=\left(0.8,0,0.6\right)$, as in the examples given in Sect. 2.2.1, are marked by solid circles and solid boxes, respectively.
Figure 14: The classical Stokes vector $\mathbf{S}_{\left(\mathrm{c}\right)}\left(\varphi\right)=\left(S_{1\left(\mathrm{c}\right)}\left(\varphi\right),S_{2\left(\mathrm{c}\right)}\left(\varphi\right),S_{3\left(\mathrm{c}\right)}\left(\varphi\right)\right)$ plotted against the phase angle $\varphi$. Its initial values $\mathbf{S}\left(\varphi=0\right)=\left(1,0,0\right)$ and $\mathbf{S}\left(\varphi\approx-1.16\times 10^{-6}\,\mathrm{rad}\right)=\left(0.8,0,0.6\right)$ are marked by solid circles and solid boxes, respectively.
@@ Appendix B Approximate analytical solutions to evolution equations
Substituting Eq. (30) into Eqs. (23)-(25), the evolution equations can be reduced as follows:
@@
First, we solve Eq. (79) for $S_{3}\left(s\right)$, and then using this solution, obtain $S_{1}\left(s\right)$ and $S_{2}\left(s\right)$, by integrating Eqs. (77) and (78), respectively:
@@
Then $S_{\mathrm{o}}$, $C$ and $\delta$ are determined by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with Eqs. (80)-(82) evaluated at $s=0$.